OpenStudy (anonymous):

x+y+z=10 5x+2y+3z=28 3x+3y+z=20

OpenStudy (anonymous):

matrice? or lots of algebra substitution

OpenStudy (anonymous):

try elimination...=)

OpenStudy (anonymous):

until you arrived at its echelon form..=)

OpenStudy (anonymous):

okay take equation (1) and multiply by 3 so you get 3x+3y+3z=30 subtract the third equation 3x+3y+z=20 and you get 2z=10 so z=5

OpenStudy (anonymous):

Wait how?

OpenStudy (anonymous):

This was the original question.. Mark, Jessica and Nate each downloaded music from the same website. Mark downloaded 10 songs in total consisting of pop, rock, and hip hop. Jessica downloaded five times as many pop songs, twice as many rock and three times as many hip hop songs as Mark. She downloaded 28 songs total. If Nate downloaded 20 songs total with three times as many pop songs, three times as many rock songs, and the same number of hip hop songs as Mark, how many songs of each type did Mark download?

OpenStudy (anonymous):

And these are the answers to choose from... Mark downloaded 4 pop songs, 2 rock songs and 4 hip hop songs. Mark downloaded 3 pop songs, 2 rock songs and 5 hip hop songs. Mark downloaded 2 pop songs, 5 rock songs and 3 hip hop songs. Mark downloaded 1 pop songs, 4 rock songs and 5 hip hop songs.

OpenStudy (anonymous):

algebra x=11 y=-6 z=5

OpenStudy (anonymous):

cannot have negative answers

OpenStudy (anonymous):

x+y+z=10 5x+2y(-)3z=28 3x+3y+z=20 lol I put a negative sign lol

OpenStudy (anonymous):

equation (3) gives 3(x+y) + z = 20, from eq (1) you get that x+y=10-z. from (3) & (1) you get that 3(10-z) + z = 20 <=> 30 -20 = 2z <=> z = 5 thus from (1) you get that x+y=10-5<=> x=5-y if you replace these to (2) you get that 5(5-y)+2y+3*5=28<=> y=4 from (1) you get that x=5-y=5-4=1 to confirm x+y+z=10<=>1+4+5=10

OpenStudy (anonymous):

(1) R+H+P=10 (2) 5P+2R+3H=28 (3)3P+3R+H=20 same procedure multiply equation (1) by 3 giving you 3R+3H+3P=30 now subtract (1)-(2) yielding 2H=10 so H=5 ( 5 hip hop songs)

OpenStudy (anonymous):

Go with the final answer that contains 5 Hip Hop songs